# Rivers under Noise

### v1

@@ 4 Main results
Then, by Lemma 9.1a) with $r=\infty$ and $l=0$,
@@
(5.5) $\lim_{s\to\infty}p^{+}\Big{(}s,z\frac{\sigma}{\sqrt{2s}}\Big{)}=\Phi(z):=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{z}\exp\Big{\{}-\frac{y^{2}}{2}\Big{\}}\,\text{\rm{d}}y.$
@@
(5.6) $\text{\rm{d}}Z_{A}(t)=\big{(}sZ_{A}(t)+A\big{)}\,\text{\rm{d}}t+\sigma\text{\rm{d}}W(t),\,t\geq s;\,Z_{A}(s)=x\in[-1,1],$
@@
By Lemma 9.1a), the probability $p_{A}(s,x)$ that $Z_{A}$ exits the interval $[-1,1]$ via 1 equals
(5.7) $\frac{q_{A}(x)-q_{A}(-1)}{q_{A}(1)-q_{A}(-1)}=\frac{\Phi\Big{(}\frac{\sqrt{2s}x}{\sigma}+\frac{2A}{\sigma\sqrt{2s}}\Big{)}-\Phi\Big{(}-\frac{\sqrt{2s}}{\sigma}+\frac{2A}{\sigma\sqrt{2s}}\Big{)}}{\Phi\Big{(}\frac{\sqrt{2s}}{\sigma}+\frac{2A}{\sigma\sqrt{2s}}\Big{)}-\Phi\Big{(}-\frac{\sqrt{2s}}{\sigma}+\frac{2A}{\sigma\sqrt{2s}}\Big{)}}.$
@@
(5.8) $\lim_{s\to\infty}p_{A}\Big{(}s,z\frac{\sigma}{\sqrt{2s}}\Big{)}=\Phi(z),$
@@
(5.9) $\zeta_{A}:=\limsup_{s\to\infty}\Big{(}\sup_{x\in[-1,1]}\mathbb{E}\big{[}T(s,x,A)\big{]}-s\Big{)}<\infty.$
@@
(5.10) $1-\rho(s,s-s^{-\alpha})\leq\exp\Big{\{}-\frac{1}{\sigma^{2}}s^{1-2\alpha}\big{(}1+o(1)\big{)}\Big{\}}$
@@
By Lemma 9.1a), the probability that $\overline{X}$ leaves the interval $[0,s]$ via $s$ equals
(5.12) $\frac{p(s-\gamma)}{p(s)}\geq 1-\rho(s,s-\gamma).$
@@
(5.13) $X_{A}(t)=A\int_{s}^{t}\exp\Big{\{}\frac{u^{2}}{2}-\frac{t^{2}}{2}\Big{\}}\,\text{\rm{d}}u+\sigma\exp\Big{\{}-\frac{t^{2}}{2}\Big{\}}\int_{s}^{t}\exp\Big{\{}\frac{u^{2}}{2}\Big{\}}\,\text{\rm{d}}W(u),\;t\geq s.$
@@
(5.14) $\leq\int_{t-1}^{t}\frac{u}{t-1}\exp\big{\{}\kappa u^{2}\big{\}}\text{\rm{d}}u+\int_{s}^{(t-1)\vee s}\frac{u}{s}\exp\big{\{}\kappa u^{2}\big{\}}\text{\rm{d}}u$
@@
(5.15) $A\Big{(}\frac{1}{t-1}+\frac{1}{s}\exp\big{\{}-t+\frac{1}{2}\big{\}}\Big{)}$
@@
The following corollary is not needed to prove Theorems 4.2 and 4.4 but will be needed to prove Theorem 4.6.
@@
(7.1) $n-n^{-\alpha_{0}}<\mathscr{R}(n)<n+n^{-\alpha_{0}}$
@@
(7.2) $\sum_{n\in\mathbb{N}}\mathbb{P}\big{(}\sup_{t\in[0,1]}Z_{n}(t)\geq n^{-\alpha}\big{)}<\infty$
@@
(7.4) $\leq 3n^{-\alpha_{0}}+\sigma\int_{0}^{t}\text{\rm{e}}^{-n(t-s)}\,\text{\rm{d}}W_{n}(s).$
@@
(7.5) $\widetilde{Y}_{n}(t)=-n^{-\alpha_{0}}\text{\rm{e}}^{-nt}+\sigma\int_{0}^{t}\text{\rm{e}}^{-n(t-s)}\,\text{\rm{d}}W_{n}(s).$
@@
For every $p>2$, Lemma 9.2 implies that
(7.6) $\mathbb{E}\Big{[}\sup_{0\leq t\leq 1}|R_{n}(t)|^{p}\Big{]}\leq C_{p}n^{1-\frac{p}{2}}.$
@@ 9 Appendix
(9.1) $\text{\rm{d}}Z(t)=b(Z(t))\,\text{\rm{d}}t+\sigma\,\text{\rm{d}}W(t),$
@@
Let $-\infty\leq l<c<r\leq\infty$, let $Z$ be the solution of (9.1) with initial condition $x\in(l,r)$ and define the exit time from the interval $(l,r)$ by $S:=\inf\{t\geq 0:Z(t)\notin(l,r)\}$. We define $p(r):=\lim_{y\uparrow r}p(y)$ if $r=\infty$ and similarly for $p(l)$.
@@
    (9.2) $\mathbb{P}\Big{(}\lim_{t\to S}Z(t)=l\Big{)}=1-\mathbb{P}\Big{(}\lim_{t\to S}Z(t)=r\Big{)}=\frac{p(r)-p(x)}{p(r)-p(l)}.$
    (Note that the scale function depends on $c$ but the right hand side of (9.2) doesn’t.)
@@
The following result is [3, Lemma 2.2] in which we replaced the constant $a_{p,\mu}$ by its numerical value given in the proof.
Lemma 9.2.
@@
Acknowledgements. Michael Scheutzow acknowledges financial support from the London Mathematical Society, grant No. 22203.

### v2

@@ 4 Main results
Then, by Lemma A.1a) with $r=\infty$ and $l=0$,
@@
$\lim_{s\to\infty}p^{+}\Big{(}s,z\frac{\sigma}{\sqrt{2s}}\Big{)}=\Phi(z):=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{z}\exp\Big{\{}-\frac{y^{2}}{2}\Big{\}}\,\text{\rm{d}}y.$ (5.5)
@@
$\text{\rm{d}}Z_{A}(t)=\big{(}sZ_{A}(t)+A\big{)}\,\text{\rm{d}}t+\sigma\text{\rm{d}}W(t),\,t\geq s;\,Z_{A}(s)=x\in[-1,1],$ (5.6)
@@
By Lemma A.1a), the probability $p_{A}(s,x)$ that $Z_{A}$ exits the interval $[-1,1]$ via 1 equals
$\frac{q_{A}(x)-q_{A}(-1)}{q_{A}(1)-q_{A}(-1)}=\frac{\Phi\Big{(}\frac{\sqrt{2s}x}{\sigma}+\frac{2A}{\sigma\sqrt{2s}}\Big{)}-\Phi\Big{(}-\frac{\sqrt{2s}}{\sigma}+\frac{2A}{\sigma\sqrt{2s}}\Big{)}}{\Phi\Big{(}\frac{\sqrt{2s}}{\sigma}+\frac{2A}{\sigma\sqrt{2s}}\Big{)}-\Phi\Big{(}-\frac{\sqrt{2s}}{\sigma}+\frac{2A}{\sigma\sqrt{2s}}\Big{)}}.$ (5.7)
@@
$\lim_{s\to\infty}p_{A}\Big{(}s,z\frac{\sigma}{\sqrt{2s}}\Big{)}=\Phi(z),$ (5.8)
@@
$\zeta_{A}:=\limsup_{s\to\infty}\Big{(}\sup_{x\in[-1,1]}\mathbb{E}\big{[}T(s,x,A)\big{]}-s\Big{)}<\infty.$ (5.9)
@@
$1-\rho(s,s-s^{-\alpha})\leq\exp\Big{\{}-\frac{1}{\sigma^{2}}s^{1-2\alpha}\big{(}1+o(1)\big{)}\Big{\}}$ (5.10)
@@
By Lemma A.1a), the probability that $\overline{X}$ leaves the interval $[0,s]$ via $s$ equals
$\frac{p(s-\gamma)}{p(s)}\geq 1-\rho(s,s-\gamma).$ (5.12)
@@
$X_{A}(t)=A\int_{s}^{t}\exp\Big{\{}\frac{u^{2}}{2}-\frac{t^{2}}{2}\Big{\}}\,\text{\rm{d}}u+\sigma\exp\Big{\{}-\frac{t^{2}}{2}\Big{\}}\int_{s}^{t}\exp\Big{\{}\frac{u^{2}}{2}\Big{\}}\,\text{\rm{d}}W(u),\;t\geq s.$ (5.13)
@@
$\leq\int_{t-1}^{t}\frac{u}{t-1}\exp\big{\{}\kappa u^{2}\big{\}}\text{\rm{d}}u+\int_{s}^{(t-1)\vee s}\frac{u}{s}\exp\big{\{}\kappa u^{2}\big{\}}\text{\rm{d}}u$ (5.14)
@@
$A\Big{(}\frac{1}{t-1}+\frac{1}{s}\exp\big{\{}-t+\frac{1}{2}\big{\}}\Big{)}$ (5.15)
@@
The following corollary is not needed to prove Theorems 4.5 and 4.7 but will be needed to prove Theorem 4.9.
@@
$n-n^{-\alpha_{0}}<\mathscr{R}(n)<n+n^{-\alpha_{0}}$ (7.1)
@@
$\sum_{n\in\mathbb{N}}\mathbb{P}\big{(}\sup_{t\in[0,1]}Z_{n}(t)\geq n^{-\alpha}\big{)}<\infty$ (7.2)
@@
$\leq 3n^{-\alpha_{0}}+\sigma\int_{0}^{t}\text{\rm{e}}^{-n(t-s)}\,\text{\rm{d}}W_{n}(s).$ (7.4)
@@
$\widetilde{Y}_{n}(t)=-n^{-\alpha_{0}}\text{\rm{e}}^{-nt}+\sigma\int_{0}^{t}\text{\rm{e}}^{-n(t-s)}\,\text{\rm{d}}W_{n}(s).$ (7.5)
@@
For every $p>2$, Lemma A.2 implies that
$\mathbb{E}\Big{[}\sup_{0\leq t\leq 1}|R_{n}(t)|^{p}\Big{]}\leq C_{p}n^{1-\frac{p}{2}}.$ (7.6)
@@ Appendix A Appendix
$\text{\rm{d}}Z(t)=b(Z(t))\,\text{\rm{d}}t+\sigma\,\text{\rm{d}}W(t),$ (A.1)
@@
Let $-\infty\leq l<c<r\leq\infty$, let $Z$ be the solution of (A.1) with initial condition $x\in(l,r)$ and define the exit time from the interval $(l,r)$ by $S:=\inf\{t\geq 0:Z(t)\notin(l,r)\}$. We define $p(r):=\lim_{y\uparrow r}p(y)$ if $r=\infty$ and similarly for $p(l)$.
@@
    $\mathbb{P}\Big{(}\lim_{t\to S}Z(t)=l\Big{)}=1-\mathbb{P}\Big{(}\lim_{t\to S}Z(t)=r\Big{)}=\frac{p(r)-p(x)}{p(r)-p(l)}.$ (A.2)
    (Note that the scale function depends on $c$ but the right hand side of (A.2) doesn’t.)
@@
The following result is [7, Lemma 2.2] in which we replaced the constant $a_{p,\mu}$ by its numerical value given in the proof.
Lemma A.2.
@@
[…]
Acknowledgements
Michael Scheutzow acknowledges financial support from the London Mathematical Society, grant No. 22203.